Bone has a Young’s modulus of about 1.8 × 1010 Pa . Under compression, it can withstand a stress of about 1.67 × 108 Pa before breaking. Assume that a femur (thigh bone) is 0.55 m long, and calculate the amount of compression this bone can withstand before breaking.


1
Expert's answer
2022-02-21T12:10:47-0500

Answer

Compression is given

"Compression=\frac{stress*L}{Y}\\=\frac{1.67*10^8*0.55}{1.8*10^{10}}\\=0.51cm"